OpenStudy (anonymous):

find the equation of the line through the points (6,7) (2, -1)

OpenStudy (campbell_st):

find the gradient 1st rise/run or\[m =(y _{2} -y _{1})/(x _{2} - x _{1})\] so ( -1 - 7)/(2-6) = -8/-4 =2 the gradient is m = 2 use the point gradient formula with either point... \[y - y _{1}=m(x - x _{1})\]

OpenStudy (anonymous):

that what i did but have have to get it equal to y which im having trouble doing...

OpenStudy (anonymous):

i have y-7=2 (x+6) so far...

OpenStudy (campbell_st):

add 7 to both sides... and expand the brackets

OpenStudy (anonymous):

sooo y=2(x+6+7) ? i put in y=2(x+13) and it was wrong.

OpenStudy (campbell_st):

y = 2(x+6) +7

OpenStudy (anonymous):

that is still coming up as wrong :(

OpenStudy (anonymous):

\[y=mx+b \] Use your two points (ordered pairs) to find the slope (rise over run or the difference of the Ys over the differences over the Xs). \[(-1-7)/(2-6)\] You can simplify this to 2. This gives you...\[y=2x+b\]. B is your Y-intercept (where the line crosses the Y-axis, or when X is zero). To solve for B with a known point and slope, fillin all the known variables and solve for B.\[-1=2*2+b\]\[-1=4+b\]\[b=-5\] Now you have the your Y-intercept. Your equation should look like this.\[y=2x-5\]